A man is trying to cross the river flowing at a speed of 5m/s to have least possible displacement by swimming at an angle of 143 degrees to the stream . 
The drift he suffers when he is crossing the same river in least possible time is (width of river = 1km)

Grade:11

1 Answers

Ishan
4 years ago
From first condition
Usin theta=-5
U sin 53=-5
U 4/5=-5
U=-25/4
PutiPut this in second condition
∆x=Vw * w/Vbw
∆x=5 * 1000/-25/4
∆x=4000/5=800m 
T=w/Vbw=1000/-25/4=4000/25=160s